# Compressed Image Generation with Denoising Diffusion Codebook Models

### v3

@@ B.1 Experiment Configurations
  • In our $256\times 256$ experiments we use the 256x256_diffusion_uncond.pt checkpoint from the official GitHub repository. In our $768\times 768$ and $512\times 512$ experiments we use Stable Diffusion 2.1 with the stabilityai/stable-diffusion-2-1 and stabilityai/stable-diffusion-2-1-base official checkpoints from Hugging Face, respectively. The different $K$, $M$, $C$ and $T$ values for each of the $512\times 512$ and the $768\times 768$ experiments plotted in Figs. 5 and 15 are summarized in Tab. 1.
@@
  • For PSC-D and PSC-R we use the same pre-trained ImageNet $256\times 256$ model as ours in the $256\times 256$ experiments, and the same Stable Diffusion 2.1 model in the $512\times 512$ experiments. We adopt the default hyper-parameters of the method as described by the authors (Elata et al., 2024), setting the number of measurements to $12\cdot 2^{i}$ for $i=0,\ldots,8$.
@@
  • For ILLM we use the MS-ILLM pre-trained models available in the official GitHub implementation. For the $512\times 512$ image size experiments we use msillm_quality_X, $\texttt{X}=2,3,4$. For the $768\times 768$ image size experiments we use msillm_quality_X, $\texttt{X}=2,3$ and msillm_quality_vloY, $\texttt{Y}=1,2$.
  • CRDR-D and CRDR-R are evaluated using quality factors of $\{0,1,2,3,4\}$, where CRDR-D uses $\beta=0$ and CRDR-R uses $\beta=3.84$, as recommended in the paper.
@@ B.2 Additional Evaluations
In Figs. 11, LABEL:, 12, LABEL:, 13, LABEL: and 14 we provide additional qualitative comparisons on the Kodak24 ($512\times 512$) and ImageNet ($256\times 256$) datasets. We additionally compare our method on images of size $768\times 768$, and present the results in Fig. 15. Our method is implemented as before, while using a Stable Diffusion 2.1 model trained on the appropriate image size (see Sec. B.1).
B.3 Decreasing the Bit Rate via Timestep Sub-Sampling
@@
B.4 Increasing the Bit Rate via Matching Pursuit
@@
In our experiments, the set of coefficients $\Gamma$ is a subset of $(0,1]$, containing $C=|\Gamma|$ values that are evenly spaced in this range. We pick $C=3$ and assess $M\in\{2,3,6\}$ for the latent space model experiments, and $M\in\{2,3,4,5\}$ for the pixel space model experiments.
B.5 Assessing the Effectiveness of Text Prompts in Compression using Text-to-Image Latent Space DDMs
@@
We assess this text-conditional approach on the $512\times 512$ SD 2.1 DDM, using CFG scales of $3,6$. We compare the performance of this conditional approach with that of the unconditional one we used in Sec. 5. The results in Fig. 16 show a disadvantage for using text-prompts for compression with our method.
Figure 11: Qualitative extreme image compression results. The presented images are taken from the Kodak24 dataset, cropped to $512\times 512$ pixels. Our compression scheme produces highly realistic decompressed outputs, while maintaining better fidelity to the original images compared to previous methods.
Figure 12: Qualitative image compression results. The presented images are taken from the Kodak24 dataset, cropped to $512\times 512$ pixels. Our compression scheme produces highly realistic decompressed outputs, while maintaining better fidelity to the original images compared to previous methods.
Figure 13: Qualitative image compression results. the presented images are taken from the ImageNet $256\times 256$ dataset. Compared to previous methods, our compression scheme produces higher perceptual quality and better fidelity to the original images.
Figure 14: Qualitative image compression results. the presented images are taken from the ImageNet $256\times 256$ dataset. Compared to previous methods, our compression scheme produces higher perceptual quality and better fidelity to the original images.
Figure 15: Quantitative image compression results on $768\times 768$ sized images. At higher bit rates, our method achieves the lowest (best) FID scores in both datasets while maintaining better distortion metrics compared to PerCo (SD). At extremely low bit rates, while PerCo (SD) shows marginally better FID scores, our method attains superior PSNR performance.
@@ C.5 Compressed Classifier Guidance
We compare our proposed CCG with the standard CG using the same unconditional diffusion model and time-dependent classifier trained on ImageNet $256\times 256$ (Deng et al., 2009; Dhariwal & Nichol, 2021). We compare the methods “on the same grounds” by using the same standard DDPM noise schedule and $T=1000$ diffusion steps. Our method is assessed with $K=256$ and $\tilde{K}=2$, while the standard CG is assessed with CG scales $s\in\{1,10,20\}$. The quantitative comparison in Tab. 2 shows that CCG achieves better (lower) FID and $\text{FD}_{\text{DINOv2}}$ scores. A visual comparison is provided in Fig. 23. Note that while using DDCM with standard CG does still produces compressed output images, decoding the produced bit-streams requires access to ${\mathbf{y}}$. Using DDCM with CCG instead sidesteps this limitation, as ${\mathbf{y}}$ is not needed for decompression.
Table 2: Quantitative comparison of compressed classifier guidance (CCG) and standard classifier guidance (CG) for ImageNet $256\times 256$ conditional image generation, using an unconditional DDM and a classifier for guidance. Our proposed CCG not only outperforms CG in terms of generation performance, but also automatically produces compressed image representations.
@@
FID $\mathbf{13.669}$ $31.548\mid 14.481\mid 14.921$
$\text{FD}_{\text{DINOv2}}$ $\mathbf{204.693}$ $459.42\mid 255.41\mid 248.34$
@@ C.6 Compressed Classifier-Free Guidance
We implement our method using SD 2.1 trained on $768\times 768$ images, adopting a DDPM noise schedule with $T=1000$ diffusion steps, $K=64$ fixed vectors in each codebook and $\tilde{K}\in\{2,3,4,6,9\}$. We compare against the same diffusion model with standard DDPM sampling, using $T=1000$ steps and CFG scales $w\in\{2,5,8,11\}$. The generative performance of both methods is assessed by computing the FID between 10k generated samples and MS-COCO, similarly to Sec. 4. Additionally, we evaluate the alignment between the outputs and the input text prompts using the CLIP score (Hessel et al., 2021) with the OpenAI CLIP ViT-L/14 model (Radford et al., 2021).

### v4

@@ B.1 Experiment Configurations
  • In our $\times 256$ 256 experiments we use the 256x256_diffusion_uncond.pt checkpoint from the official GitHub repository. In our $\times 768$ 768 and $\times 512$ 512 experiments we use Stable Diffusion 2.1 with the stabilityai/stable-diffusion-2-1 and stabilityai/stable-diffusion-2-1-base official checkpoints from Hugging Face, respectively. The different $K$, $M$, $C$ and $T$ values for each of the $\times 512$ 512 and the $\times 768$ 768 experiments plotted in Figs. 5 and 15 are summarized in Tab. 1.
@@
  • For PSC-D and PSC-R we use the same pre-trained ImageNet $\times 256$ 256 model as ours in the $\times 256$ 256 experiments, and the same Stable Diffusion 2.1 model in the $\times 512$ 512 experiments. We adopt the default hyper-parameters of the method as described by the authors (Elata et al., 2024), setting the number of measurements to $12\cdot 2^{i}$ for $i=0,\ldots,8$.
@@
[…]
@@ B.2 Additional Evaluations
In Figs. 11, LABEL:, 12, LABEL:, 13, LABEL: and 14 we provide additional qualitative comparisons on the Kodak24 ($\times 512$ 512) and ImageNet ($\times 256$ 256) datasets. We additionally compare our method on images of size $\times 768$ 768, and present the results in Fig. 15. Our method is implemented as before, while using a Stable Diffusion 2.1 model trained on the appropriate image size (see Sec. B.1).
B.3 Numerical Results
Tabs. 2, LABEL:, 3, LABEL:, 4, LABEL: and 5 include the numerical results that appear in Fig. 5. Additionally, Tabs. 6, LABEL: and 7 include the numerical results for Fig. 15.
B.4 Decreasing the Bit Rate via Timestep Sub-Sampling
@@
B.5 Increasing the Bit Rate via Matching Pursuit
@@
In our experiments, the set of coefficients $\Gamma$ is a subset of (0,1](0,1]( 0 , 1 ], containing $C=|\Gamma|$ values that are evenly spaced in this range. We pick $C=3$ and assess $M\in\{2,3,6\}$ for the latent space model experiments, and $M\in\{2,3,4,5\}$ for the pixel space model experiments.
B.6 Assessing the Effectiveness of Text Prompts in Compression using Text-to-Image Latent Space DDMs
@@
We assess this text-conditional approach on the $\times 512$ 512 SD 2.1 DDM, using CFG scales of 3,63,63 , 6. We compare the performance of this conditional approach with that of the unconditional one we used in Sec. 5. The results in Fig. 16 show a disadvantage for using text-prompts for compression with our method.
Figure 11: Qualitative extreme image compression results. The presented images are taken from the Kodak24 dataset, cropped to $\times 512$ 512 pixels. Our compression scheme produces highly realistic decompressed outputs, while maintaining better fidelity to the original images compared to previous methods.
Figure 12: Qualitative image compression results. The presented images are taken from the Kodak24 dataset, cropped to $\times 512$ 512 pixels. Our compression scheme produces highly realistic decompressed outputs, while maintaining better fidelity to the original images compared to previous methods.
Figure 13: Qualitative image compression results. the presented images are taken from the ImageNet $\times 256$ 256 dataset. Compared to previous methods, our compression scheme produces higher perceptual quality and better fidelity to the original images.
Figure 14: Qualitative image compression results. the presented images are taken from the ImageNet $\times 256$ 256 dataset. Compared to previous methods, our compression scheme produces higher perceptual quality and better fidelity to the original images.
Figure 15: Quantitative image compression results on $\times 768$ 768 sized images. At higher bit rates, our method achieves the lowest (best) FID scores in both datasets while maintaining better distortion metrics compared to PerCo (SD). At extremely low bit rates, while PerCo (SD) shows marginally better FID scores, our method attains superior PSNR performance.
@@
Table 2: Compression quantitative evaluations, for the ImageNet $\times 256$ 256 dataset.
Method BPP FID LPIPS PSNR
BPG 0.056 144.322 0.460 23.771
0.118 111.821 0.330 26.085
0.248 81.555 0.218 28.329
0.503 60.512 0.127 30.864
0.941 42.714 0.064 33.602
HiFiC 0.220 42.610 0.055 27.713
0.395 31.567 0.033 29.849
0.584 25.125 0.021 31.622
IPIC 0.095 45.367 0.211 24.470
0.148 38.413 0.152 26.011
0.228 36.301 0.132 27.664
0.550 27.075 0.074 31.178
0.945 20.689 0.044 33.800
PSC-D 0.019 147.317 0.591 21.648
0.039 127.275 0.496 23.572
0.077 105.225 0.404 25.605
0.155 86.787 0.316 27.711
PSC-P 0.019 67.457 0.431 18.575
0.039 59.119 0.353 20.178
0.077 51.464 0.268 22.649
0.155 44.589 0.194 25.107
Ours 0.018 68.339 0.419 19.624
0.023 63.731 0.365 20.566
0.032 57.027 0.296 21.877
0.041 52.157 0.251 22.795
0.053 47.834 0.220 23.401
0.069 43.972 0.185 24.199
0.091 40.787 0.165 24.643
0.122 36.890 0.134 25.587
0.183 31.977 0.102 26.765
0.381 28.247 0.076 27.949
0.594 25.705 0.063 28.694
0.808 24.165 0.055 29.211
1.021 22.938 0.050 29.596
Table 3: Compression quantitative evaluations, for the Kodak24 $\times 512$ 512 dataset.
Method BPP FID LPIPS PSNR
BPG 0.037 148.306 0.540 24.211
0.044 141.575 0.513 24.638
0.062 128.095 0.461 25.539
0.088 114.256 0.407 26.492
0.121 104.289 0.357 27.375
0.167 92.423 0.307 28.346
0.227 82.184 0.260 29.359
CRDR-D 0.114 94.475 0.257 27.988
0.210 76.015 0.167 30.069
CRDR-R 0.114 35.676 0.094 27.315
0.210 28.391 0.057 29.373
HiFiC 0.191 31.713 0.068 27.376
0.363 25.734 0.042 29.466
ILLM 0.085 36.080 0.110 25.683
0.159 28.556 0.072 27.296
0.304 24.448 0.044 29.395
PSC-P 0.025 47.334 0.473 18.516
0.049 40.691 0.371 20.080
0.095 35.314 0.266 21.683
PerCo (SD) 0.033 37.019 0.307 19.017
0.127 26.418 0.145 22.325
Ours 0.030 32.031 0.222 22.066
0.038 29.117 0.190 22.551
0.050 25.647 0.161 23.013
0.095 24.215 0.138 23.606
0.149 23.199 0.124 24.069
0.309 22.260 0.108 24.665
SD 2.1
Encoder-Decoder
Bound
0.071 26.428
Table 4: Compression quantitative evaluations, for the CLIC2020 $\times 512$ 512 dataset.
Method BPP FID LPIPS PSNR
BPG 0.040 112.453 0.451 24.777
0.046 106.936 0.428 25.213
0.064 96.809 0.380 26.149
0.088 86.578 0.331 27.138
0.119 77.203 0.285 28.072
0.160 68.827 0.242 29.049
0.214 60.534 0.204 30.063
CRDR-D 0.106 52.459 0.193 28.934
0.188 40.228 0.124 31.062
0.355 27.793 0.070 33.705
CRDR-R 0.106 14.292 0.074 28.352
0.188 10.022 0.045 30.452
0.355 6.225 0.025 33.028
HiFiC 0.173 14.000 0.053 28.878
0.322 10.157 0.032 31.040
ILLM 0.081 12.666 0.087 26.898
0.146 9.923 0.056 28.662
0.271 8.246 0.034 30.840
0.391 7.222 0.024 32.233
PSC-P 0.025 15.007 0.435 18.153
0.049 13.527 0.335 20.110
0.095 11.879 0.229 22.276
PerCo (SD) 0.033 13.896 0.287 18.111
0.127 7.888 0.128 22.453
Ours 0.030 9.459 0.186 22.630
0.038 8.386 0.160 23.171
0.050 7.755 0.137 23.748
0.095 7.340 0.116 24.472
0.149 7.164 0.103 25.008
0.309 6.825 0.088 25.782
SD 2.1
Encoder-Decoder
Bound
0.056 27.901
Table 5: Compression quantitative evaluations, for the DIV2K $\times 512$ 512 dataset.
Method BPP FID LPIPS PSNR
BPG 0.058 152.623 0.490 22.793
0.068 146.219 0.463 23.217
0.096 134.029 0.407 24.124
0.134 122.895 0.348 25.095
0.327 87.879 0.192 28.049
CRDR-D 0.152 89.766 0.203 26.618
0.274 69.466 0.118 28.762
CRDR-R 0.152 39.291 0.081 26.110
0.274 31.422 0.048 28.275
HiFiC 0.226 39.832 0.066 26.138
ILLM 0.104 42.591 0.107 24.579
0.187 34.103 0.069 26.206
0.332 29.267 0.043 28.229
PSC-P 0.025 48.350 0.474 15.869
0.049 44.267 0.386 17.470
0.095 38.406 0.281 19.220
PerCo (SD) 0.033 40.944 0.325 16.518
0.127 29.516 0.155 20.483
Ours 0.030 36.283 0.238 19.769
0.038 32.917 0.206 20.319
0.050 30.724 0.177 20.881
0.095 29.322 0.150 21.566
0.149 28.011 0.132 22.107
0.309 26.756 0.114 22.867
SD 2.1
Encoder-Decoder
Bound
0.076 24.878
Table 6: Compression quantitative evaluations, for the CLIC2020 $\times 768$ 768 dataset.
Method BPP FID LPIPS PSNR
BPG 0.032 115.009 0.451 25.547
0.038 109.587 0.429 25.984
0.054 99.482 0.385 26.925
0.075 89.013 0.338 27.915
0.102 79.456 0.296 28.840
0.138 71.107 0.255 29.811
0.185 62.617 0.219 30.812
HiFiC 0.154 9.359 0.052 29.745
ILLM 0.006 62.327 0.447 21.304
0.009 41.479 0.340 22.703
0.072 7.852 0.084 27.808
0.134 5.966 0.055 29.567
PerCo (SD) 0.003 30.409 0.517 15.339
0.032 12.869 0.269 19.018
0.126 5.419 0.122 23.387
Ours 0.007 23.862 0.404 19.672
0.008 19.521 0.354 20.532
0.010 15.559 0.314 21.207
0.014 11.362 0.262 22.116
0.017 8.722 0.227 22.722
0.022 6.753 0.192 23.366
0.042 5.051 0.156 24.136
0.066 4.549 0.133 24.739
0.137 4.132 0.108 25.650
SD 2.1
Encoder-Decoder
Bound
0.055 28.850
Table 7: Compression quantitative evaluations, for the DIV2K $\times 768$ 768 dataset.
Method BPP FID LPIPS PSNR
BPG 0.049 139.204 0.485 23.504
0.058 132.393 0.460 23.928
0.081 120.682 0.406 24.841
0.114 107.472 0.350 25.826
HiFiC 0.205 23.077 0.063 27.096
ILLM 0.007 92.242 0.485 19.234
0.011 71.777 0.369 20.522
0.093 24.253 0.102 25.446
0.169 18.386 0.067 27.085
PerCo (SD) 0.003 51.915 0.519 14.436
0.032 31.378 0.311 17.364
0.126 16.487 0.148 21.317
Ours 0.007 48.315 0.469 17.194
0.008 42.932 0.422 17.858
0.010 37.680 0.383 18.413
0.014 31.038 0.328 19.246
0.017 26.599 0.286 19.832
0.022 22.073 0.247 20.432
0.042 18.330 0.201 21.167
0.066 16.396 0.172 21.780
0.137 14.662 0.138 22.707
SD 2.1
Encoder-Decoder
Bound
0.071 25.919
@@ C.5 Compressed Classifier Guidance
We compare our proposed CCG with the standard CG using the same unconditional diffusion model and time-dependent classifier trained on ImageNet $\times 256$ 256 (Deng et al., 2009; Dhariwal & Nichol, 2021). We compare the methods “on the same grounds” by using the same standard DDPM noise schedule and $T=1000$ diffusion steps. Our method is assessed with $K=256$ and $\tilde{K}=2$, while the standard CG is assessed with CG scales $s\in\{1,10,20\}$. The quantitative comparison in Tab. 8 shows that CCG achieves better (lower) FID and $\text{FD}_{\text{DINOv2}}$ scores. A visual comparison is provided in Fig. 23. Note that while using DDCM with standard CG does still produces compressed output images, decoding the produced bit-streams requires access to ${\mathbf{y}}$. Using DDCM with CCG instead sidesteps this limitation, as ${\mathbf{y}}$ is not needed for decompression.
Table 8: Quantitative comparison of compressed classifier guidance (CCG) and standard classifier guidance (CG) for ImageNet $\times 256$ 256 conditional image generation, using an unconditional DDM and a classifier for guidance. Our proposed CCG not only outperforms CG in terms of generation performance, but also automatically produces compressed image representations.
@@
FID $\mathbf{13.669}$ 31.54814.48114.92131.548$\mid$ 14.481$\mid$ 14.92131.548 ∣ 14.481 ∣ 14.921
$\text{FD}_{\text{DINOv2}}$ $\mathbf{204.693}$ 459.42255.41248.34459.42$\mid$ 255.41$\mid$ 248.34459.42 ∣ 255.41 ∣ 248.34
@@ C.6 Compressed Classifier-Free Guidance
We implement our method using SD 2.1 trained on $\times 768$ 768 images, adopting a DDPM noise schedule with $T=1000$ diffusion steps, $K=64$ fixed vectors in each codebook and $\tilde{K}\in\{2,3,4,6,9\}$. We compare against the same diffusion model with standard DDPM sampling, using $T=1000$ steps and CFG scales $w\in\{2,5,8,11\}$. The generative performance of both methods is assessed by computing the FID between 10k generated samples and MS-COCO, similarly to Sec. 4. Additionally, we evaluate the alignment between the outputs and the input text prompts using the CLIP score (Hessel et al., 2021) with the OpenAI CLIP ViT-L/14 model (Radford et al., 2021).